40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 8 13:20:07 2011, steve, Update, PEM, vac envelope at atm for 36 days vent70.jpg
    Reply  Fri Jun 8 15:24:08 2012, steve, Update, VAC, RGA scan at day 276 and pressure plots 40m71pdM276d.pngpd71m276d.pngpd71d276.png
Message ID: 5366     Entry time: Thu Sep 8 13:20:07 2011     Reply to this: 6790
Author: steve 
Type: Update 
Category: PEM 
Subject: vac envelope at atm for 36 days 

We tried not to open chambers above 10,000 particles of 0.5 micron cf/min

New items going in:                       2 rasor beam traps, 5 badly oxidized old silver plated  setscrew with spring loaded tips......to be replaced in the future, viton tips for eq screws....some are lose, gold plated allen wrench installed at ITMX bottom, reglued magnet on ITMX

Bad hardware things found:          nylon ball "locking elements" on OSEM locking set screws with screwdriver  slot, lose  1064 nm filter on OSEM pd

Attachment 1: vent70.jpg  90 kB  | Hide | Hide all
vent70.jpg
ELOG V3.1.3-